[4 – (-3) +1] / [2 – 2 / (-2)]
Answer by dolly(163)

[4 – (-3) +1] / [2 – 2 / (-2)]
= (4+3+1)/(2+1) [as -2/-2 = +1]
= 8/3
= 2 2/3